# I’M HOI: Inertia-aware Monocular Capture of 3D Human-Object Interactions

### v1

@@ B.3 Dataset Statistics
[…]
@@ C.1 More Results
In former sections, we show the robustness of I’m-HOI under severe occlusions. Here, we further exhibit sequential capture results of I’m-HOI to demonstrate spatial-temporal coherence. Figure 10 illustrates that our approach captures accurate and consistent human-object spatial arrangements in a time window, which validates that our proposed network learns reasonable interaction distributions and recognizes continuous interaction behaviors from input with hybrid modality.
@@ C.2 More Comparisons
We also present more qualitative comparisons on sequential capture results with baselines in Figure 11 and Figure 12. It can be observed that even within an extremely short time interval (around 0.07 seconds), jittery object tracking results occur in image-based baselines [104, 85] which model the static interactions but ignore temporal information. While, video-based method [86] gives temporally consistent but erroneous predictions without inertial measurements, especially for the tracking of object rotational motion. In stark contrast, I’m-HOI makes use of both visual cues and IMU signals, cooperating with design of object-oriented mesh alignment feedback and category-specific interaction prior, which benefits to consistent and correct results.
@@ C.3 Ablation on Regularization Terms
To further evaluate the effectiveness of regularization terms in interaction diffusion filter training, we compare the performance of full model with downgraded versions that discard each term. As reported in Table 6, $\mathcal{L}_{\text{off}}$, restricting object to a more specific and precise region, improves accuracy. $\mathcal{L}_{\text{consist}}$ regularizes predicted joint rotations to comply with detected 3D joints after forward kinematics, which prevents overfitting to pseudo ground-truth annotations. Both $\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ improves performance in temporal dimension. However, only applying $\mathcal{L}_{\text{vel}}$ can lead to oversmooth results since the physical dynamics is lost. Incorporating second-order supervision $\mathcal{L}_{\text{imu}}$ is verified profitable for not only smooth but also physically plausible capturing.
[…]
Figure 11: More qualitative comparisons. I’m-HOI performs better than baselines on sequential data.
Figure 12: More qualitative comparisons. I’m-HOI performs better than baselines on sequential data.

### v2

@@ B.3 Dataset Statistics
We present a comprehensive overview of the contents of IMHD2 in Table 5. It reveals that, for each object, we curated a wide array of interaction patterns involving different human body segments. Complementary to existing datasets characterized by numerous participants, extensive recording frames and super dense views, Figure 9 illustrates that IMHD2 offers a more challenging, diverse and quality collection of motion data focusing on object-oriented interactions. Specifically, measured through metrics such as average motion velocity and jitter, IMHD2 encompasses more dynamic interaction motions with better smoothness. Moreover, IMU data is concurrently collected alongside RGB images, serving not only to align with ground-truth annotations, but also as network input to enhance accuracy and efficiency in motion capture.
@@ C.1 More Results
In preceding sections, we have demonstrated the robustness of I’m-HOI under severe occlusions. Expanding on this, we now present sequential capture results of I’m-HOI to showcase spatial-temporal coherence. Figure 10 illustrates that our approach captures accurate and consistent human-object spatial arrangements within a temporal context, which validates that our proposed network learns reasonable interaction distributions and recognizes continuous interaction behaviors from input data featuring a hybrid modality.
@@ C.2 More Comparisons
We also present additional qualitative comparisons of sequential capture results with baselines in Figure 11 and Figure 12. It can be observed that even within an extremely short time interval (approximately 0.07 seconds), the image-based baselines [104, 85] exhibit jittery object tracking results, focusing on static interactions while disregarding temporal information. Conversely, the video-based method [86] yields temporally consistent but erroneous predictions without inertial measurements, particularly evident in tracking object rotational motions. In stark contrast, I’m-HOI makes use of both visual cues and IMU signals, cooperating with the design of object-oriented mesh alignment feedback and category-specific interaction prior. This combination contributes significantly to achieving consistent and correct results.
@@ C.3 Ablation on Regularization Terms
To further evaluate the effectiveness of the regularization terms in training of interaction diffusion filter, we conduct a comparative analysis of the full model against downgraded versions that exclude individual terms. As reported in Table 6, the inclusion of $\mathcal{L}_{\text{off}}$ restricts objects to a more specific and precise region. $\mathcal{L}_{\text{consist}}$ enforces predicted joint rotations to align with detected 3D joints after forward kinematics, which prevents overfitting to pseudo ground-truth annotations. Both $\mathcal{L}_{\text{vel}}$ and $\mathcal{L}_{\text{imu}}$ contribute to improve performance in the temporal domain. However, only applying $\mathcal{L}_{\text{vel}}$ may lead to oversmooth results due to the loss of physical dynamics. Incorporating second-order supervision $\mathcal{L}_{\text{imu}}$ is verified beneficial, not only for smooth results but also for capturing physically plausible interaction motions.
Figure 10: Additional qualitative results of I’m-HOI on IMHD2. We present sequential RGB images, captured motion from camera view and top-view visualizations.
Figure 11: Additional qualitative comparisons. I’m-HOI outperforms baselines on sequential data.
Figure 12: Additional qualitative comparisons. I’m-HOI outperforms baselines on sequential data.